OpenStudy (anonymous):

I have read and read this and I guess I am dumb, but I do not understand this at all....could someone please help me? Evaluate a polynomial for a given value of the variable. Demonstrate the process with an example.

OpenStudy (anonymous):

I think the question is asking to make an polynomial yourself and solve it when x equals an arbitrary number.

OpenStudy (anonymous):

I am not sure how to do this, everything that I have read, does not make sense to me, Not sure how to reply to it....

OpenStudy (anonymous):

f(x) = x^3 + 2x - 3 let x= 0 f(0) = x^3 + 2^0 -3 = -3

OpenStudy (anonymous):

A polynomial can be quadratic, cubic and so on. For example, a quadratic is y=x^2. Solve for y when x equals 1.

OpenStudy (anonymous):

Thank you so much I am def going to have my work cut out for me in this class.